2.Find the probability that a randomly selected senior high school student spends less than 21 hours or greater than 30 hours.


1
Expert's answer
2021-05-07T09:11:42-0400

Mean is 24 hours with a standard deviation of 4 hours.


Let's find the opposite (the probability that a student spends more than 21 and less than 30 hours), than subtract this probability from 1, and that will be the answer.


The probability that 21 < X < 30 is equal to the blue area under the curve




Since "\mu" = 24 and "\sigma =4"  we have:

P( 21 < X < 30 ) = P(21-24 < X - "\mu" < 30 - 24) = P"(\frac{21-24}{4}<\frac{X-\mu}{\mu} <\frac{30-24}{4} )"

Since "Z = \frac{X-\mu}{\sigma}", "\frac{21-24}{4} = -0.75" and "\frac{30-24}{4} = 1.5" we have:

P( 21 < X < 30) = P(-0.75 < Z < 1.5)

Use the standard normal table to conclude that:

P(-0.75 < Z < 1.5) = 0.7066

So, P' = 1 - 0.7066 = 0.2934

Answer: P = 0.2934.


P.S. Here is the z-score table: